OpenStudy (anonymous):

Solve the equation: 3(2x - 4) = 4(x + 1)

OpenStudy (anonymous):

3(2x - 4) = 4(x + 1) 6x - 12 = 4x + 4 6x - 4x - 12= 4x - 4x + 4 2x - 12 = + 4 2x - 12 + 12 = 4 + 12 2x = 16 2x/2 = 16/2 x = 8